# Approaching ballistic motion in 3D simulations of gamma-ray burst jets in realistic binary neutron star merger environments

### v1

@@ Abstract
Aims. The outcome of relativistic (magneto)hydrodynamic simulations of jets piercing through post-merger environments is often used as input to compute afterglow signals to be compared with observations. However, for reliable comparisons, the jet propagation should be followed until nearly ballistic regimes, in which the jet acceleration is essentially over and the angular structure is no longer evolving. This condition is typically reached in 2D simulations, but not in 3D. Our goal is to extend a (specific) jet simulation in 3D up to a nearly ballistic phase, analysing the overall dynamical evolution from the jet breakout.
Methods. Our work is based on a previous 3D magnetohydrodynamic jet simulation employing a realistic environment imported from a BNS merger simulation, extended here far beyond the evolution time originally covered. After approximately 3 seconds of the jet evolution on the original spherical grid, we remap the system into a uniform Cartesian grid and reach about 10 seconds without loss of resolution.
Results. The specific jet considered here struggles to pierce the dense surroundings, resulting in a rather asymmetrical emerging outflow with relatively low Lorentz factor. The analysis of the energy conversion processes and corresponding acceleration shows that at the end of our simulation 98% of the energy is in kinetic form. Moreover, at that time the angular structure is frozen. We thus obtain suitable inputs for computing the afterglow emission. Our procedure is general and applicable to any jet simulation of the same kind.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are bright explosive events, characterized by an early prompt emission phase detected in the sub-MeV band, followed by a shallower, radio-to-TeV afterglow emission phase that can last for up to days, months, or even years after the initial burst. Numerous investigations indicate that both the prompt and afterglow emission arise from the energy dissipation of a collimated relativistic outflow (‘jet’), launched from a compact central engine (e.g. a hyper-accreting black hole or a highly magnetized, fast-spinning neutron star). The jet propagates through the surrounding medium generating shocks and triggering various dynamical effects (including turbulence and magnetic reconnection) leading to radiative dissipation through the synchrotron and inverse-Compton processes (e.g. Kumar & Zhang, 2015). These investigations further suggest that the progenitor system responsible for the formation of the central engine powering the jet is either a core-collapsing massive star or a merging compact binary system (e.g. Zhang, 2018). In particular, the association between at least a fraction of ‘short’ GRBs (sGRBs; defined as those with prompt emission phase lasting less than 2 s) and binary neutron star (BNS) mergers has been recently confirmed by the concomitant observation of gravitational waves (GWs) from a coalescing BNS (event named GW170817) and a short duration burst of gamma-rays (GRB 170817A), with the latter detected $\mathrm{\simeq\!1.74}$ s after the estimated time of merger (Abbott et al., 2017b, c, a). Analyses of the afterglow emission accompanying GRB 170817A have further revealed that a collimated relativistic jet, launched from the merger remnant and propagating along a direction misaligned with respect to Earth, is responsible for the electromagnetic (EM) emission, imprinting its angular structure and energetics into the observed signals (Mooley et al., 2018a, b; Lazzati et al., 2018; Ghirlanda et al., 2019). Providing conclusive evidence that BNS mergers can launch collimated relativistic outflows and power GRBs, such a discovery has promoted unprecedented efforts in exploring this connection (e.g. Ciolfi, 2020a; Margutti & Chornock, 2021).
@@
The angular structure of a sGRB jet from BNS coalescence is shaped during the launch and the subsequent propagation through the post-merger environment surrounding the compact central engine (Salafia & Ghirlanda, 2022, and refs. therein). After breaking out of such environment, the jet angular structure is expected to still evolve for a while until it eventually freezes out. At the same time, the jet keeps accelerating until essentially all the available energy is in kinetic form. When those conditions are reached (ballistic regime), the jet expands freely in the interstellar medium (ISM) on orders of magnitude longer timescales, until the interaction with the latter starts having relevant dynamical effects and the afterglow radiation is produced.
@@
The afterglow signal is therefore expected to carry key information about the jet launching and the following propagation that is complementary to what can be inferred from GW signals. In order to exploit such information, it is however necessary to accurately model the jet propagation dynamics and, in this context, numerical simulations plausibly represent the most promising approach.
@@
Since the GRB 170817A event, the evolution of sGRB jets from BNS mergers has been extensively investigated through relativistic hydrodynamic or magnetohydrodynamic simulations, performed in two or three spatial dimensions (e.g. Lazzati et al., 2017; Murguia-Berthier et al., 2021; Geng et al., 2019; Kathirgamaraju et al., 2019; Nathanail et al., 2020, 2021; Urrutia et al., 2021, 2023; Gottlieb et al., 2020, 2021, 2022; Gottlieb & Nakar, 2022; Gottlieb et al., 2023; García-García et al., 2023). While most of these simulations employ hand-made surrounding environments to study the propagation of incipient jets, the first investigations based on more realistic environments directly imported from the outcome of BNS merger simulations were recently accomplished, both in absence and in presence of magnetic fields (Pavan et al., 2021, 2023; Lazzati et al., 2021).111see also Nativi et al. (2021, 2022)
@@
The outcome of these simulations, in terms of jet energetics and angular structure, can be in principle used as input to compute the corresponding multi-wavelength afterglow signals, depending on the viewing angle, the density of the ISM, and other parameters (e.g. Lazzati et al. 2018). However, the obtained light curves are reliable only if the jet expansion has reached a nearly ballistic regime. Such a condition is typically met in studies based on two-dimensional (2D) simulations (e.g. Lazzati et al. 2018; Xie et al. 2018; Urrutia et al. 2021), but not in the case of three-dimensional (3D) simulations (e.g. Kathirgamaraju et al. 2019; Nathanail et al. 2021; Nativi et al. 2022).
@@
In this work, we present the methods and procedure that we developed to extend, without loss of resolution, simulations like the one presented in Pavan et al. (2023, P23 henceforth) up to a nearly ballistic jet expansion phase. Taking the simulation of P23 as a reference model, we show how the evolution can be extended up to $\approx\!10$ s after jet launching, when its angular structure is frozen and its energy is $\approx\!98$% kinetic. Besides demonstrating the methods, we study the dynamics and energetics of the specific jet examined, and extract the final angular profiles of energy and Lorentz factor that can be employed as input to compute the corresponding afterglow light curves.
@@
The paper is structured as follows. In Section 2, we describe the numerical setup of the reference jet simulation, along with the methods implemented to achieve the extended evolution up to a nearly ballistic phase and to define the jet ‘head’ for the following analysis. In Section 3, we present our results in terms of jet dynamical evolution and energetics. In Section 4, we perform a detailed study of the jet structure and the evolving angular distribution of energy and Lorentz factor. In Section 5, we estimate the saturation Lorentz factor and deceleration radius for the jet model at hand. Finally, in Section 6, we summarize the present work and comment on future perspectives.
@@ 2.1 Reference Model
In the following, we introduce and briefly describe the input model employed in this work, corresponding to the fiducial case presented in P23.
@@
The post-merger environment surrounding the central remnant, through which the incipient jet has to propagate, is directly imported, in terms of rest-mass density, pressure, velocity, and magnetic field, from the outcome of a general-relativistic magnetohydrodynamic simulation of a BNS merger presented in Ciolfi (2020b). This realistic, magnetized environment is interpolated onto a new grid in 3D spherical coordinates, to be evolved with the special-relativistic magnetohydrodynamic module of the PLUTO code (Mignone et al., 2007). A 90 axis rotation is also applied so that the new polar axis is orthogonal to the orbital axis of the merging BNS (along which the jet is then injected), avoiding any effect of the polar axis on the jet evolution. An excision radius of 380 km is applied in the new domain to cut out the central region, so that general relativity effects that are not accounted for in PLUTO (and are not relevant for the jet evolution at large scales) can be safely neglected. Logarithmic spacing in the radial direction is imposed to obtain the highest resolution closer to the inner boundary. The outer radial boundary is set to $r_{\mathrm{max}}$ = $2.5\times 10^{6}$ km. Grid spacing along $r,\theta,\phi$, at $r\!=\!380$ km, is $\approx 4.4,4.4$, and $4.7$ km, respectively, corresponding to a resolution that is at least as high as in the original merger simulation (at that distance from the origin). Moreover, the uniform density and pressure characterizing the artificial atmosphere in the merger simulation is redefined into a variable floor scaling as $r^{-6.5}$.
@@
For the time evolution, a piecewise parabolic reconstruction, the HLL Riemann solver, and the third order Runge-Kutta time stepping are adopted. To handle the divergence-free condition of the magnetic field, the Hyperbolic Divergence Cleaning technique is employed (see, e.g. Mignone & Tzeferacos 2010). Furthermore, we use the Taub-Matthews equation of state (an analytical approximation of the one derived by Synge in 1957), implemented in the PLUTO code (see Mignone et al. 2005; Mignone & McKinney 2007). We also note that the (Newtonian) gravitational pull from the central object is included, using the mass of the BNS merger remnant, namely 2.596 M.
@@
After the data import, the post-merger environment is evolved in PLUTO employing radial boundary conditions at the excision radius that are set to reproduce the angle-averaged time evolution of the different physical quantities as directly extrapolated from the BNS merger simulation. A substitution method developed in P23 is used to exploit in full the data from the merger simulation up to the latest available time ( 255 ms after merger). Then, the environment is further evolved for another 100 ms. At 355 ms after merger, the merger remnant is assumed to collapse to a black hole (BH). For a time window of 30 ms, the effects of the collapse are accounted for by introducing a fading acceleration term in the equations of motion aimed at reproducing the gradual decrease of radial pressure gradients due to the BH formation and subsequent accretion. Finally, at $t_{jet}$ = 385 ms after merger, a relativistic jet is launched. In particular, a magnetized, uniformly rotating, axisymmetric jet is manually injected with the following parameters: half-opening angle $\Theta_{\mathrm{j}}=10^{\circ}$, initial Lorentz factor $\Gamma_{\mathrm{j}}=3$, terminal Lorentz factor $\Gamma_{\infty}=300$, and initial two-sided luminosity of $L_{j}=3\times 10^{51}$ erg/s that fades away exponentially with a characteristic timescale of 0.3 s. The density and pressure angular profiles within the jet are computed by solving a transverse balance equation between total pressure gradient, centrifugal force, and magnetic tension. The jet magnetic field profile is the same adopted by Geng et al. (2019) (see also Martí, 2015). The level of magnetization in the jet is described by the parameter
$\frac{L_{\mathrm{j}}-L_{\mathrm{j},\text{HD}}}{L_{\mathrm{j}}}\simeq 8\leavevmode\nobreak\ \text{\%}$ (1)
@@ 2.3 Jet head definition
Figure 2: Kinetic power computed at different radii for the three representative times. In each panel, the blue, red, and green dot-dashed lines correspond to 10, 15, and 20% of the maximum $dE/dt$, respectively. The continuous magenta line marks the initial luminosity of the jet at injection (north side only).
@@ 3.1 Jet dynamics
Figure 3: Meridional view of rest-mass density, Lorentz factor, kinetic energy density (left to right) at the jet breakout time ($t_{BO}\!\simeq\!350$ ms). The white lines give an indicative separation between the region containing the jet and the surrounding material.
@@
From the distribution of the Lorentz factor, we observe that the high velocity part of the outflow consists of the radially expanding jet front followed by a short tail, surrounded by a spherical shell accelerated at earlier times which is however of extremely low density and energetically unimportant (see below). The observed radial stratification in the tail is associated to the fading behaviour of the injected luminosity. During the evolution, there is no significant radial spread of the high velocity region ($\Gamma\gtrsim 2$), while the jet continues to expand laterally, becoming more and more similar to a portion of a spherical shell with a nearly constant half-opening angle of about 30. At this stage, the jet is no longer interacting with the environment surrounding the merger site and only undergoes further acceleration and slight morphological evolution while expanding across the artificial floor medium, which has negligible impact on the dynamics.
In the first seconds of evolution the jet entrains a large amount of environmental material (see Fig. 1) which is only in part imported in the new Cartesian domain. During our simulation the rest-mass density distribution, shown in the second column of Fig. 4, indicates that such material is expanding slowly, with a low density funnel previously excavated by the jet along the $y$-axis. At the lower boundary ($y=10^{5}$ km), we see the effects of the exponentially decaying velocity $v_{y}$ imposed to damp further injection of matter and energy, as discussed in the previous Section 2.2. Analogous effects can be seen in the pressure and magnetic field strength (third and fourth columns).
To get a better visualization of the distribution of the outflow material in the whole domain, a three-dimensional plot of the kinetic energy density at the end of the simulation is reported in Fig. 5. It is evident that the energy distribution is highly non-axisymmetric (see also Sect. 4) and the contribution of the shell observed in the Lorentz factor (see Figs. 1 and 4) is negligible at large angles ($\gtrsim\!30^{\circ}$) from the jet injection axis.
Figure 5: 3D rendering of the kinetic energy density at the final simulation time ($\!\approx\!9.5$ s after jet injection). Higher transparency is applied to the lowest values of energy to allow for a better visualization. Spatial scale is limited to $y\in[1,3.5]\times 10^{6}$ km.

### v2

@@ Abstract
Aims. The outcome of relativistic (magneto)hydrodynamic simulations of jets piercing through post-merger environments is often used as input to compute afterglow signals that can be compared with observations. However, for reliable comparisons, the jet propagation should be followed until nearly ballistic regimes, in which the jet acceleration is essentially over and the angular structure is no longer evolving. This condition is typically reached in 2D simulations, but not in 3D ones. Our goal is to extend a (specific) jet simulation in 3D up to a nearly ballistic phase and analyse the overall dynamical evolution from the jet breakout.
Methods. Our work is based on a previous 3D magnetohydrodynamic jet simulation employing a realistic environment imported from a BNS merger simulation, extended here far beyond the evolution time originally covered. After approximately 3 seconds of the jet evolution on the original spherical grid, we remapped the system into a uniform Cartesian grid and reached about 10 seconds without loss of resolution.
Results. The specific jet considered here struggled to pierce the dense surroundings, resulting in a rather asymmetrical emerging outflow with a relatively low Lorentz factor. Analysis of the energy conversion processes and corresponding acceleration showed that at the end of our simulation, 98% of the energy is in kinetic form. Moreover, at that time the angular structure is frozen. We thus obtained suitable inputs for computing the afterglow emission. Our procedure is general and applicable to any jet simulation of the same kind.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are bright explosive events characterised by an early prompt emission phase detected in the sub-MeV band followed by a shallower radio-to-TeV afterglow emission phase that can last for up to days, months, or even years after the initial burst. Numerous investigations have indicated that both the prompt and afterglow emission arise from the energy dissipation of a collimated relativistic outflow (‘jet’) launched from a compact central engine (e.g. a hyper-accreting black hole or a highly magnetised fast-spinning neutron star). The jet propagates through the surrounding medium, generating shocks and triggering various dynamical effects (including turbulence and magnetic reconnection) and leading to radiative dissipation through the synchrotron and inverse-Compton processes (e.g. Kumar & Zhang, 2015). These investigations have further suggested that the progenitor system responsible for the formation of the central engine powering the jet is either a core-collapsing massive star or a merging compact binary system (e.g. Zhang, 2018). In particular, the association between at least a fraction of ‘short’ GRBs (sGRBs; defined as those with a prompt emission phase lasting less than 2 s) and binary neutron star (BNS) mergers has been recently confirmed by the concomitant observation of gravitational waves (GWs) from a coalescing BNS (event named GW170817) and a short duration burst of gamma rays (GRB 170817A), with the latter detected $\mathrm{\simeq\!1.74}$ s after the estimated time of merger (Abbott et al., 2017b, c, a). Analyses of the afterglow emission accompanying GRB 170817A have further revealed that a collimated relativistic jet, launched from the merger remnant and propagating along a direction misaligned with respect to Earth, is responsible for the electromagnetic (EM) emission, and it has imprinted its angular structure and energetics onto the observed signals (Mooley et al., 2018a, b; Lazzati et al., 2018; Ghirlanda et al., 2019). Providing conclusive evidence that BNS mergers can launch collimated relativistic outflows and power GRBs, such a discovery has promoted unprecedented efforts in exploring this connection (e.g. Ciolfi, 2020a; Margutti & Chornock, 2021).
@@
The angular structure of a sGRB jet from BNS coalescence is shaped during the launch and the subsequent propagation through the post-merger environment surrounding the compact central engine (Salafia & Ghirlanda, 2022; and refs. therein). After breaking out of such an environment, the jet angular structure is expected to still evolve for a while until it eventually freezes out. At the same time, the jet keeps accelerating until essentially all the available energy is in kinetic form. When those conditions are reached (ballistic regime), the jet expands freely in the interstellar medium (ISM) on orders of magnitude longer time scales until the interaction with the latter starts having relevant dynamical effects and the afterglow radiation is produced.
@@
The afterglow signal is therefore expected to carry key information about the jet launching and the following propagation that is complementary to what can be inferred from GW signals. In order to exploit such information, it is necessary to accurately model the jet propagation dynamics, and in this context, numerical simulations plausibly represent the most promising approach.
@@
Since the GRB 170817A event, the evolution of sGRB jets from BNS mergers has been extensively investigated through relativistic hydrodynamic or magnetohydrodynamic simulations, performed in two or three spatial dimensions (e.g. Lazzati et al., 2017; Murguia-Berthier et al., 2021; Geng et al., 2019; Kathirgamaraju et al., 2019; Nathanail et al., 2020, 2021; Urrutia et al., 2021, 2023; Gottlieb et al., 2020, 2021, 2022; Gottlieb & Nakar, 2022; Gottlieb et al., 2023; García-García et al., 2023). While most of these simulations employ hand-made surrounding environments to study the propagation of incipient jets, the first investigations based on more realistic environments directly imported from the outcome of BNS merger simulations were recently accomplished, both in the absence and in presence of magnetic fields (Pavan et al., 2021, 2023; Lazzati et al., 2021).111see also Nativi et al. (2021, 2022)
@@
The outcome of these simulations, in terms of jet energetics and angular structure, can in principle be used as input to compute the corresponding multi-wavelength afterglow signals, depending on the viewing angle; the density of the ISM; and other parameters (e.g. Lazzati et al. 2018). However, the obtained light curves are reliable only if the jet expansion has reached a nearly ballistic regime. Such a condition is typically met in studies based on two-dimensional simulations (e.g. Lazzati et al. 2018; Xie et al. 2018; Urrutia et al. 2021) but not in the case of three-dimensional simulations (e.g. Kathirgamaraju et al. 2019; Nathanail et al. 2021; Nativi et al. 2022).
@@
In this work, we present the methods and procedure that we developed to extend, without loss of resolution, simulations such as the one presented in Pavan et al. (2023; P23 henceforth) up to a nearly ballistic jet expansion phase. Taking the simulation of P23 as a reference model, we show how the evolution can be extended up to $\approx\!10$ s after jet launching, when its angular structure is frozen and its energy is $\approx\!98$% kinetic. Besides demonstrating the methods, we study the dynamics and energetics of the specific jet examined, and we extract the final angular profiles of energy and Lorentz factor that can be employed as input to compute the corresponding afterglow light curves.
@@
The paper is structured as follows. In Section 2, we describe the numerical setup of the reference jet simulation, along with the methods implemented to achieve the extended evolution up to a nearly ballistic phase and to define the jet ‘head’ for the following analysis. In Section 3, we present our results in terms of jet dynamical evolution and energetics. In Section 4, we perform a detailed study of the jet structure and the evolving angular distribution of energy and Lorentz factor. In Section 5, we estimate the saturation Lorentz factor and deceleration radius for the jet model at hand. Finally, in Section 6, we summarise the present work and comment on future perspectives.
@@ 2.1 Reference model
In the following, we introduce and briefly describe the input model employed in this work, which corresponds to the fiducial case presented in P23. The post-merger environment surrounding the central remnant through which the incipient jet has to propagate was directly imported in terms of rest-mass density, pressure, velocity, and magnetic field from the outcome of a general-relativistic magnetohydrodynamic simulation of a BNS merger presented in Ciolfi (2020b). This realistic magnetised environment was interpolated onto a new grid in 3D spherical coordinates in order to be evolved with the special-relativistic magnetohydrodynamic module of the PLUTO code (Mignone et al., 2007). A 90 axis rotation was also applied so that the new polar axis was orthogonal to the orbital axis of the merging BNS (along which the jet was then injected), avoiding any effect of the polar axis on the jet evolution. An excision radius of 380 km was applied in the new domain to cut out the central region so that general relativity effects that are not accounted for in PLUTO (and are not relevant for the jet evolution at large scales) can be safely neglected. Logarithmic spacing in the radial direction was imposed to obtain the highest resolution closer to the inner boundary. The outer radial boundary was set to $r_{\mathrm{max}}$ = $2.5\times 10^{6}$ km. Grid spacing along $r,\theta,\phi$, at $r\!=\!380$ km, was $\approx 4.4,4.4$, and $4.7$ km, respectively, corresponding to a resolution that is at least as high as in the original merger simulation (at that distance from the origin). Moreover, the uniform density and pressure characterising the artificial atmosphere in the merger simulation was redefined into a variable floor scaling as $r^{-6.5}$.
@@
For the time evolution, a piecewise parabolic reconstruction, the HLL Riemann solver, and the third order Runge-Kutta time stepping were adopted. To handle the divergence-free condition of the magnetic field, the Hyperbolic Divergence Cleaning technique was employed (see e.g. Mignone & Tzeferacos 2010). Furthermore, we used the Taub-Matthews equation of state (an analytical approximation of the one derived by Synge in 1957), implemented in the PLUTO code (see Mignone et al. 2005; Mignone & McKinney 2007). We also note that the (Newtonian) gravitational pull from the central object was included, using the mass of the BNS merger remnant, namely 2.596 M.
@@
After the data import, the post-merger environment was evolved in PLUTO employing radial boundary conditions at the excision radius that were set to reproduce the angle-averaged time evolution of the different physical quantities as directly extrapolated from the BNS merger simulation. A substitution method developed in P23 was used to exploit in full the data from the merger simulation up to the latest available time ( 255 ms after merger). Then, the environment was further evolved for another 100 ms. At 355 ms after merger, the merger remnant was assumed to collapse to a black hole (BH). For a time window of 30 ms, the effects of the collapse were accounted for by introducing a fading acceleration term in the equations of motion aimed at reproducing the gradual decrease of radial pressure gradients due to the BH formation and subsequent accretion. Finally, at $t_{jet}$ = 385 ms after merger, a relativistic jet was launched. In particular, a magnetised uniformly rotating axisymmetric jet was manually injected with the following parameters: half-opening angle $\Theta_{\mathrm{j}}=10^{\circ}$, initial Lorentz factor $\Gamma_{\mathrm{j}}=3$, terminal Lorentz factor $\Gamma_{\infty}=300$, and initial two-sided luminosity of $L_{j}=3\times 10^{51}$ erg/s that fades away exponentially with a characteristic time scale of 0.3 s. The density and pressure angular profiles within the jet were computed by solving a transverse balance equation between total pressure gradient, centrifugal force, and magnetic tension. The jet magnetic field profile is the same as what was adopted by Geng et al. (2019) (see also Martí, 2015). The level of magnetisation in the jet is described by the parameter
$\frac{L_{\mathrm{j}}-L_{\mathrm{j},\text{HD}}}{L_{\mathrm{j}}}\simeq 8\leavevmode\nobreak\ \text{\%},$ (1)
@@ 2.3 Jet head definition
Figure 2: Kinetic power computed at different radii for the three representative times. In each panel, the blue, red, and green dot-dashed lines correspond to 10%, 15%, and 20% of the maximum $dE/dt$, respectively. The continuous magenta line marks the initial luminosity of the jet at injection (north side only).
@@ 3.1 Jet dynamics
Figure 3: Meridional view of rest-mass density, Lorentz factor, and kinetic energy density (left to right) at the jet breakout time ($t_{BO}\!\simeq\!350$ ms). The white lines give an indicative separation between the region containing the jet and the surrounding material.
@@
From the distribution of the Lorentz factor, we observe that the high velocity part of the outflow consists of the radially expanding jet front followed by a short tail surrounded by a spherical shell accelerated at earlier times, which is of extremely low density and energetically unimportant (see below). The observed radial stratification in the tail is associated to the fading behaviour of the injected luminosity. During the evolution, there is no significant radial spread of the high velocity region ($\Gamma\gtrsim 2$), while the jet continues to expand laterally, becoming more and more similar to a portion of a spherical shell with a nearly constant half-opening angle of about 30. At this stage, the jet is no longer interacting with the environment surrounding the merger site and only undergoes further acceleration and slight morphological evolution while expanding across the artificial floor medium, which has negligible impact on the dynamics.
In the first seconds of evolution, the jet entrains a large amount of environmental material (see Fig. 1) that is only in part imported in the new Cartesian domain. During our simulation the rest-mass density distribution, shown in the second column of Fig. 4, indicates that such material is expanding slowly, with a low density funnel previously excavated by the jet along the $y$-axis. At the lower boundary ($y=10^{5}$ km), we see the effects of the exponentially decaying velocity $v_{y}$ imposed to damp further injection of matter and energy, as discussed in the previous Section 2.2. Analogous effects can be seen in the pressure and magnetic field strength (third and fourth columns).
To get a better visualisation of the distribution of the outflow material in the whole domain, a three-dimensional plot of the kinetic energy density at the end of the simulation is reported in Fig. 5. It is evident that the energy distribution is highly non-axisymmetric (see also Sect. 4) and the contribution of the shell observed in the Lorentz factor (see Figs. 1 and 4) is negligible at large angles ($\gtrsim\!30^{\circ}$) from the jet injection axis.
Figure 5: Three-dimensional rendering of the kinetic energy density at the final simulation time ($\!\approx\!9.5$ s after jet injection). Higher transparency is applied to the lowest values of energy to allow for a better visualisation. Spatial scale is limited to $y\in[1,3.5]\times 10^{6}$ km.